OpenStudy (anonymous):

simplify the rational expression 3-3x/3x^2+6x-9 help s:

OpenStudy (anonymous):

someone?

OpenStudy (shamil98):

Factor the denominator and numerator..

OpenStudy (shamil98):

\[\huge \frac{ 3-3x }{ 3x^2+6x-9 } = \frac{ 3(1-x) }{ 3(x^2+2x-3) }\] cancel the threes to get your reduced form..

OpenStudy (shamil98):

He isn't asking for the minimum, bashirk, and screenshotting wolfram really doesn't help..

OpenStudy (anonymous):

thanks @shamil98

OpenStudy (anonymous):

Oh. Sry thn